OpenStudy (anonymous):

What is the solution to the equation? 2-z/3=12 A.z=-42 B.z=-34 C.z=-30 D.z=42 @iGreen

OpenStudy (igreen):

Is this it? \(2 - \dfrac{z}{3} = 12\)

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Or is it this one? \(\dfrac{2 - z}{3} = 12\)

OpenStudy (anonymous):

the first one

OpenStudy (igreen):

Okay..so first we subtract 2 to both sides.. what's 12 - 2?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Can you subtract that? @Alpha

OpenStudy (igreen):

@Alpha_Female02

OpenStudy (anonymous):

10

OpenStudy (igreen):

Yes, that gives us: \(- \dfrac{z}{3} = 10\) Now we multiply 3 * 10.

OpenStudy (anonymous):

=30

OpenStudy (igreen):

\(-z = 30\) Yes, now we divide -1 to both sides, what's 30 / -1?

OpenStudy (anonymous):

-30

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (anonymous):

wait isn't that one of the answer choices?

OpenStudy (igreen):

Yes..that's C..-30.